a. tahri found that the composition of a compound is magnesium 20.20%, sulfur 26.60%, and oxygen 53.20% by chemical analysis. calculate its empirical formula.
b. if the molecular mass is 120g/mol what is the molecular formula
Answer #1

a)

we have mass of each elements as:

Mg: 20.2 g

S: 26.6 g

O: 53.2 g

Divide by molar mass to get number of moles of each:

Mg: 20.2/24.31 = 0.8309

S: 26.6/32.07 = 0.8294

O: 53.2/16.0 = 3.325

Divide by smallest to get simplest whole number ratio:

Mg: 0.8309/0.8294 = 1

S: 0.8294/0.8294 = 1

O: 3.325/0.8294 = 4

  

So empirical formula is:MgSO4

Answer: MgSO4

b)

Molar mass of MgSO4,

MM = 1*MM(Mg) + 1*MM(S) + 4*MM(O)

= 1*24.31 + 1*32.07 + 4*16.0

= 120.38 g/mol

Now we have:

Molar mass = 120.0 g/mol

Empirical formula mass = 120.38 g/mol

Multiplying factor = molar mass / empirical formula mass

= 120.0/120.38

= 1

So molecular formula is:MgSO4

Answer: MgSO4